OpenStudy (anonymous):

Solve the equation for y. 7x - 3y = 4

OpenStudy (lgbasallote):

what's the problem here?

OpenStudy (anonymous):

right there ^

OpenStudy (anonymous):

\[7x-3y=4\]\[-3y=-7x+4\]\[y=\left( -7x+4 \right)\div-3\]Is that what you're looking for?

OpenStudy (lgbasallote):

i meant where are you finding it difficult?

OpenStudy (anonymous):

one of theseee VVV A. y = x + B. y = 7x - 4 C. y = x + D. y = x -

OpenStudy (anonymous):

like i dont know how to do it .

OpenStudy (anonymous):

Not sure if you're missing parts to those answers, but it looks like its neither of them to me. :/

OpenStudy (anonymous):

dammm .